OpenStudy (anonymous):

Where does the graph of y = 2x2 + 5x − 3 cross the x-axis?

OpenStudy (anonymous):

first you got to foil it out so it would be (2x-1)(x+3)

OpenStudy (anonymous):

then you have to put each parentheses equal to zero then solve for x

OpenStudy (australopithecus):

the x-intercept (when the graph crosses the x-axis) is when y = 0 so you will get points that look like this, (x,0) so set y = 0 and solve for x 0 = 2x^2 + 5x − 3

OpenStudy (australopithecus):

you can factor like mayb40 did or use the quadratic formula

OpenStudy (australopithecus):

0=(2x-1)(x+3) so figure out how to make (2x-1) = 0 2x = 1 x = (1/2) thus when x =(1/2), y = 0 0=(2(1/2)-1)((1/2)+3) 0 = (1-1)((1/2)+3) 0 = 0((1/2) + 3) 0 = 0 now just do the same thing for x+3 and you have your answer